OpenStudy (anonymous):

A circle is represented by the equation shown below. (x + 3)2 + (x - 2)2 = 4 Which statement is true? Answer The circle is centered at (3, 2) and has a radius of 2. The circle is centered at (-3, 2) and has a diameter of 2. The circle is centered at (-3, 2) and has a radius of 2. The circle is centered at (-3, -2) and has a diameter of 2.

Parth (parthkohli):

@lalaly :)

OpenStudy (lalaly):

Do it do it do it :D

Parth (parthkohli):

Thank you:)

Parth (parthkohli):

See, the equation of a circle looks like: \(\Large \color{MidnightBlue}{\Rightarrow (x - b)^2 + (y - h)^2 = r^2 }\) (h,b) are center co-ordinates. r is the radius.

Parth (parthkohli):

The equation is same as: \(\Large \color{MidnightBlue}{\Rightarrow (x - (-3))^2 + (y - 2)^2 = 4 }\) Center coordinates are (-3,2), radius is sqrt4 = 2. Diameter is two times radius, so it is 4.

Parth (parthkohli):

I mean diameter is 4 and radius is 2.

OpenStudy (anonymous):

Okay, this is still confusing, but I think I am starting t understand it now.

Parth (parthkohli):

Yeah.. @lalaly you're the best :D continue

OpenStudy (lalaly):

nothing is confusing about it @CoffeeEyes just write the general form everytime u see the equation of the circle, and compare it to the equation you have